OpenStudy (anonymous):

Which set of integers is a Pythagorean triple? A. 2, 3, 6 B. 3, 4, 5 C. 4, 5, 9 D. 5, 7, 12

OpenStudy (dan815):

so a^2+b^2=c^2 must be satified

OpenStudy (dan815):

plug in those numbers and see which ones work

OpenStudy (dan815):

ill do the first one for you, 2^2+3^2=13 which is not equal to 6^2 or 36 so the first option of 2,3,6 being a Pythagorean triple is wrong

OpenStudy (dan815):

try the 2nd one yourself

OpenStudy (anonymous):

So it has to come up to 36?

OpenStudy (dan815):

no the 3 numbers must satisfy the equation that is all

OpenStudy (dan815):

x^2+y^2=z^2 or a^2+b^2=c^2 what ever you want to call the variables

OpenStudy (dan815):

put in the set of 3 numbers you are given into the eqution and see which one works

OpenStudy (dan815):

the largest number for c and the other 2 numbers for a and b

OpenStudy (dan815):

For example ill give you a pythagorean triple 5,12,13 is a pythagorean triple because 5^2+12^2=13^2<--- this is a true statement, calculator ---------------------------------------------------- 2^2+3^2=6^2<---- False Statement 3^2+4^2=5^2<----- Is this True or False?